( b ) At5 har pressure having unit Iniil5s of fluid withi specific volumne of 0 . 20 ) ini " / kg , contained in essel espands to 0 . 6 bar according pvc where n2 nut ce is constarit . Calculate work done during the process

thermodymamics
Answer #1

Note : As the problem question is not clear. I are giving a possible answers for what understood. J.Pf. [ Isothermal processna If [Polytrope Process pon-c nr 2. P,= 5x105 N/m Pg = 0.6*105 N/m Vi = 0.2 m/kg &xo.29 - 12 ton Y2 = 0.57m3kg. Wook done =